Draw the ellipse with the equation: 2𝑥 2 + 𝑦 2 + 4𝑥 − 9𝑦 = 7


1
Expert's answer
2021-02-24T12:18:35-0500

first of all, we need to transform the equation to the canonical form

"\boxed{\frac {(x-x_0)^2} {a^2} + \frac{(y-y_0)^2} {b^2} = 1}" - "x_0" and "y_0" - center coordinates


"2x^2 + 4x + y^2-9y=7"


"(2x^2 +4x +2) +(y^2 -9y+20.25) = 29.25"


"2(x+1)^2+(y-4.5)^2 = 29.25"


"\dfrac{(x+1)^2}{14.625} + \dfrac{(y - 4.5)^2}{29.25} = 1"


O(-1; 4.5) - center

"a = \sqrt{14.625} \thickapprox 3.8" - minor semiaxis

"b = \sqrt{29.25} \thickapprox 5.4" - major semiaxes